The focal length ____________ when the refractive index of the lens is increased and __________ when the curvature radius of the lens is increased.
increases / increases
decreases / increases
increases / decreases
decreases / decreases
Answer #1

Relationship between focal length and refractive index is, \(\frac{1}{f}=(n-1)\left(\frac{1}{R_{1}}-\frac{1}{R_{2}}\right)\)

Here, \(f\) is focal length, \(n\) is refractive index, and \(R_{1}, R_{2}\) are radii of curvature. From the above relation, it is clear that \(\mathrm{f}\) and \(\mathrm{n}\) are inversely proportional to each other. Focal length and radius of curvature are directly proportional to each other.

Thus, the correct answer is decreases/increases.